Lee spends his entertainment budget on video games and movies. Video games cost $60

each and movies cost $20. If Lee bought an extra video game (keeping movies constant) he

would gain 30 utils (i.e. unit of utilities), if he watched an extra movie (keeping video games

constant) he would gain 15 utils. From this information we can conclude that Lee:

a) should buy fewer video games and go see more movies

b) should buy more video games and go see fewer movies.

c) should buy more video games and see more movies

d) should not change anything
Answer #1

Answer

Option a

a) should buy fewer video games and go see more movies

A consumer maximizes utility by consuming at marginal utility per dollar from both good is equal.

MU per dollar =MU/Price

MU per dollar of video game =30/60=0.5

MU per dollar of movie =15/20

=0.75

the MU per dollar from movies is higher than the video games so the consumer should increase the consumption of movies and reduce the consumption of video games.